OpenStudy (anonymous):

Find the zeros of the polynomial function and state the multiplicity of each. f(x)=5(x+7)^2(x-7)^3

OpenStudy (anonymous):

ok this is the vertex, Not a Conic Section

OpenStudy (anonymous):

my answer is only if your doing a vertex

jimthompson5910 (jim_thompson5910):

hint: if x-k is a factor, then k is a zero

jimthompson5910 (jim_thompson5910):

example (x-2)(x+3) = 0 x-2=0 or x+3=0 x=2 or x = -3

OpenStudy (anonymous):

OOOOOOOH i new that,

jimthompson5910 (jim_thompson5910):

the exponent over each factor tells you the multiplicity of that factor

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

-7, multiplicity 2; 7, multiplicity 3 4, multiplicity 1; -7, multiplicity 3; 7, multiplicity 3 -7, multiplicity 3; 7, multiplicity 2 4, multiplicity 1; 7, multiplicity 1; -7, multiplicity 1 those are my answers to choose from

jimthompson5910 (jim_thompson5910):

using my example, can you determine some of the roots?

OpenStudy (primeralph):